# Model-free reinforcement learning with noisy actions for automated experimental control in optics

### v2

@@ C.7 Algorithms
We tested six different algorithms with their standard hyperparameters in StableBaselines3 with the parameters in Tables 1 and 2 for $P_{\text{goal}}=0.8,0.85,0.9$ for either $10^{5}$ (for $P_{\text{goal}}=0.8,0.85$) steps or $5\cdot 10^{5}$ (for $P_{\text{goal}}=0.9$) training steps. The results are shown in Figure 11 Panel (b)-(d).
Figure 11: Comparison of different maximum actions and algorithms. The plots show the normalized return against the training step for different maximum actions $a_{\text{max}}$ and TQC in (a) or for six different algorithms: TQC, SAC, TD3, PPO, DDPG and A2C in (b)-(d). (b) shows this for $P_{\text{goal}}=0.8$, (c) for $P_{\text{goal}}=0.85$, and (b) for $P_{\text{goal}}=0.9$. The other parameters are chosen as in Tables 1 and 2. The plots show the mean with $2\sigma$ error bands created by multiple runs and smoothing.
We can see that in the first $10^{5}$ steps, A2C and PPO always perform worst, and DDPG is next in line. However in Figure 11 (d), we can see that for $P_{\text{goal}}=0.9$ PPO catches up to DDPG around training step $1.5\cdot 10^{5}$. SAC, TQC, and TD3 perform much better than these three. TD3 is nearly always worse than SAC. TQC always has a drop in the middle region but catches up to SAC in the end. Overall, SAC seems to be the best algorithm for this task when used on the virtual testbed. In contrast to that, in our physical experiments, TQC slightly outperforms SAC for $P_{\text{goal}}=0.85$.
@@ C.8 Effect of noise on the learning process
[…]
@@ D.1 Different goal powers
We run experiments with $P_{\text{goal}}=0.85,0.86,0.87,0.88,0.9$ using TQC. The normalized return is shown in Figure 13 (a). We can see that, just like in the virtual testbed, the training needs longer to converge the higher the goal. It is interesting that there is a big gap between the agents with $P_{\text{goal}}=0.87$ and $P_{\text{goal}}=0.88$. Please note that we performed these training runs (except for $P_{\text{goal}}=0.85$ only once and draw our conclusions from there.
@@ D.3 Pre-training on virtual testbed
Furthermore, we want to know if pretraining on the virtual testbed helps with the experiment’s training times. We tested both an agent pre-trained on the virtual testbed without noise and on a version of the virtual testbed with noise. As noise, we sample random values for each of the actuators from the dead-zone characterization in Figure 1 (c) and reduce the absolute value of the action by these sampled values.
@@ Appendix E Algorithm Hyperparameters
We use the default hyperparameters in StableBaselines3 (incl. contrib), Version 2.3.0 [34] or the way they appear in their tutorials. For completeness, we list them here and print the ones that are not default but used in the tutorial in bold.
@@
    learning rate: 0.0003, replay buffer size: 1000000, learning starts after 100 steps, batch size: 256, soft update coefficient: 0.005, discount factor: 0.99, update model every step, do 1 gradient step after each rollout, no added action noise, update target network every 1 step,

### v3

@@ C.7 Algorithms
We tested seven different algorithms with their standard hyperparameters in StableBaselines3 with the parameters in Tables 1 and 2 for $P_{\text{goal}}=0.8,0.85,0.9$ for either $10^{5}$ (for $P_{\text{goal}}=0.8,0.85$) steps or $5\cdot 10^{5}$ (for $P_{\text{goal}}=0.9$) training steps. The results are shown in Figure 11 Panel (b)-(d).
Figure 11: Comparison of different maximum actions and algorithms. The plots show the normalized return against the training step for different maximum actions $a_{\text{max}}$ and TQC in (a) or for seven different algorithms: TQC, SAC, TD3, PPO, DDPG, A2C, and CrossQ in (b)-(d). (b) shows this for $P_{\text{goal}}=0.8$, (c) for $P_{\text{goal}}=0.85$, and (b) for $P_{\text{goal}}=0.9$. The other parameters are chosen as in Tables 1 and 2. The plots show the mean with $2\sigma$ error bands created by multiple runs and smoothing.
We can see that in the first $10^{5}$ steps, A2C and PPO always perform worst, and DDPG is next in line. However in Figure 11 (d), we can see that for $_{\text{goal}}=0.9 PP$O catches up to DDPG around training step $1.5\cdot 10^{5}$. CrossQ, SAC, TQC, and TD3 perform much better than these three. TD3 is nearly always worse than SAC. TQC always has a drop in the middle region but catches up to SAC in the end. CrossQ outperforms the other algorithms for $P_{\text{goal}}=0.9$, while for lower goals, its behavior is similar to the other well-performing algorithms. Overall, CrossQ, followed by SAC, seems to be the best algorithm for this task when used on the virtual testbed. TQC suffers from a drop in performance in the middle of training compared to SAC. In contrast to that, in our physical experiments, TQC slightly outperforms SAC for $P_{\text{goal}}=0.85$.
@@ C.8 Effect of noise on the learning process
As expected, the return for higher noise levels is generally smaller than the one for no noise, but for each of the presented noise levels, the agents are still able to learn. For $P_{\text{goal}}=0.85$, the graphs for $\mathcal{N}=0,\ 1$ are very similar and only for the higher noise factors ($\mathcal{N}=2,\ 3$) the learning curve clearly differs. That means that for a moderate goal power, the noise in the experiment does not affect the agent as much. However, this is different for $P_{\text{goal}}=0.9$, the graph for $\mathcal{N}=1$ is grouped with the ones for $\mathcal{N}=2,\ 3$. Hence, the agents’ learning curves are significantly impacted by the noise level found in the experiment.
@@ D.1 Different goal powers
We run experiments with $P_{\text{goal}}=0.85,0.86,0.87,0.88,0.9$ using TQC. The normalized return is shown in Figure 13 (a). We can see that, just like in the virtual testbed, the training needs longer to converge the higher the goal. It is interesting that there is a big gap between the agents with $P_{\text{goal}}=0.87$ and $P_{\text{goal}}=0.88$. Please note that we performed these training runs (except for $P_{\text{goal}}=0.85$) only once and draw our conclusions from there.
@@ D.3 Pre-training on virtual testbed
Furthermore, we want to know if pre-training on the virtual testbed helps with the experiment’s training times. We tested both an agent pre-trained on the virtual testbed without noise and on a version of the virtual testbed with noise. As noise, we sample random values for each of the actuators from the dead-zone characterization in Figure 1 (c) and reduce the absolute value of the action by these sampled values.
@@ Appendix E Algorithm Hyperparameters
We use the default hyperparameters in StableBaselines3 (incl. contrib), Version 2.3.0 (Raffin et al., 2021) or the way they appear in their tutorials. For completeness, we list them here and print the ones that are not default but used in the tutorial in bold.
@@
    learning rate: 0.0003, replay buffer size: 1000000, learning starts after 100 steps, batch size: 256, soft update coefficient: 0.005, discount factor: 0.99, update model every step, do 1 gradient step after each rollout, no added action noise, update target network every 1 step
  • CrossQ
    learning rate: 0.0003, replay buffer size: 1000000, learning starts after 100 steps, batch size: 256, discount factor: 0.99, update model every step, do 1 gradient step after each rollout, no added action noise